### v1

# When Should Selfish Miners Double-Spend?

Abstract
Although, both double-spending and selfish-mining attacks have been extensively studied since the “Bitcoin” whitepaper of Nakamoto and the “majority is not enough” paper of Eyal and Sirer, there has been no rigorous stochastic analysis of an attack that combines the two, except for the complicated MDP models. In this paper, we first combine stubborn and selfish mining attacks, i.e., construct a strategy where the attacker acts stubborn until its private branch reaches a certain length and then switches to act selfish. We provide the optimal stubbornness for each parameter regime. Next, we provide the maximum stubbornness that is still more profitable than honest mining and argue a connection between the level of stubbornness and the $k$-confirmation rule. We show that, at each attack cycle, if the level of stubbornness is higher than $k$, there is a risk of double-spending which comes at no-cost to the adversary. The result can be seen as a guide for picking $k$ in the $k$-confirmation rule in a blockchain design. At each cycle, for a given stubbornness level, we rigorously formulate how great the risk of double-spending is. We provide the minimum double-spend value needed for an attack to be profitable in the regimes where the scheme is less profitable than honest mining. We further modify the attack in the stubborn regime in order to conceal the attack and increase the double-spending probability. Finally, we evaluate the results and provide the optimal and the maximum stubbornness levels for each parameter regime as well as the revenue. As a case study, with Bitcoin’s $k=6$ block confirmation rule, we evaluate the revenue and double-spending risk of the attacks for each pool parameter.
1 Introduction
After Nakamoto’s seminal paper[1] that laid the foundations of blockchains aiming for large-scale consensus, there have been numerous works studying the related security guarantees and flaws [2]. The private attack, investigated initially in a non-rigorous manner in the Bitcoin whitepaper [1], aims to replace a transaction in the blockchain ledger with a conflicting transaction, allowing the attacker to double-spend its funds. Rosenfeld [3] showed rigorously that the private attack has exponentially decaying probability of success with time under the longest chain protocol. More specifically, as time passes, more public blocks are mined on top of a transaction that is being attacked, and it becomes less likely for a competing private chain to overtake the longest chain and undo the transaction in question.
The celebrated selfish mining attack of Eyal and Sirer [4] exposed another vulnerability of the longest chain protocol by showing that it is not incentive-compatible. More specifically, [4] showed that a pool of miners, who constitute $\alpha$ proportion of the total hashrate in the network, can increase their revenue ratio $\rho$ by keeping the blocks they have mined private for a time according to the selfish mining algorithm, resulting in $\rho>\alpha$. This algorithm, analyzed via the use of a Markov chain, aims to waste the resources of non-pool miners by replacing their blocks with the private blocks of the pool and hence stealing their revenues. Moreover, the extra revenue of selfish mining, i.e., $\rho-\alpha$, grows even further as the pool size increases, which attracts the rational miners to join the pool, eventually overtaking the chain completely.
Later on, block withholding attacks and their effects on the revenue have gained further attention and been studied extensively [5, 6, 7, 8]. Sapirshtein et. al. [6] use Markov Decision Processes (MDP) to provide $\epsilon$-optimal policies for selfish mining attacks that result in more profit than the algorithm of [4]. In a concurrent work, Nayak et. al. [7] provide an in-depth look at variations of selfish mining attacks, including the-now-famous stubborn mining, and analyze their revenues using Monte Carlo simulation techniques. Further, Gervais et. al. [8] create an extended MDP problem where the model takes into account additional parameters such as the confirmation rule, network propagation dynamics, mining costs and eclipse attack. The authors of [8] further modify the model to consider the profitability of the double-spending attack via the MDP and investigate the impacts of the parameters on the optimal policies. Similar to the study of [8], [9] builds an MDP problem to investigate the profitability of the double-spending attack by slightly modifying the MDP model of [6] that replaces the rewards of override blocks by double-spend blocks defined therein. The authors in [9] further modify the MDP to investigate the strategies that optimally combine selfish mining attacks with double-spend attacks assuming a constant relation between the profits from the two attacks. The authors of [10] investigate a novel efficient approach to solve average reward ratio MDP of [6] using probabilistic termination methods. Other generalizations of the MDP model also exist that analyze selfish mining in different proof systems and protocols [11, 12, 13, 14, 15]. Another line of research [16, 17] investigates the profitability of double-spending attacks using stochastic methods.
An important feature of selfish mining attack is the network influence of the pool, which determines the proportion of non-pool miners who favor the adversarial chain in case of fork-ties. This parameter, which is called $\gamma$, provides a threshold for the profitability of selfish mining, i.e., the minimum pool size needed for the selfish mining attack to be profitable. In other words, $\gamma$ determines the minimum $\alpha$ needed such that $\rho(\alpha,\gamma)>\alpha$. The authors of [4] argue that the current bitcoin protocol, which picks the first chain seen by a miner in case of a tie, is vulnerable as the cost of increasing network influence in this case is minimal for the adversary. This in turn, resulted in search for new consensus policies that increase the threshold and the security [4, 18, 19, 20]. Other studies provide methods to effectively calculate network influence $\gamma$ under various network models [21, 22]. We refer the interested reader to the survey in [23] for a further read on selfish mining and double-spending attacks and related defense mechanisms.
1.1 Our Contributions
In this paper, for the first time in the literature, we provide a rigorous and explicit stochastic analysis of the combination of the two celebrated attacks: double-spending and selfish mining. To do so, we define a selfish mining attack strategy, that not only tracks the length difference between the pool’s private chain and the public chain but also its length. Tracking the length of the private chain allows us to determine whether the attack can result in double-spending or not in addition to the revenue steal from the honest miners. Under the $k$-confirmation rule, we argue that if the private branch of a selfish miner reaches length $k+1$, then the double-spending comes at no-cost. Thus, the question becomes whether it is profitable in terms of mining revenues to keep the private branch until it reaches length $k+1$. In other words, we investigate whether the new mining strategy is profitable, which is essentially a mix of selfish mining and stubborn mining that we call $L$-stubborn mining that acts stubborn until the private chain reaches length $L$, then acts selfish, i.e., releases the chain when the difference between the private and the public branches reduces to $1$.
For each $\alpha$ and $\gamma$, we find the optimal $L^{*}$, i.e., the best strategy, under the condition that the adversary does not do any trailing behind, i.e., it accepts the honest branch whenever it falls behind. Similarly, we find largest $L$, that we call $\bar{L}$, such that $\rho>\alpha$, i.e., maximum stubbornness that still results in higher revenues than honest mining. Next, we argue that if $\bar{L}>k$, then double-spending comes at no-cost. Hence, the results not only give a new understanding of selfish and stubborn mining but also provide parameter regimes of $\alpha$ and $\gamma$ where transactions are at risk from the perspective of selfish mining. It is already known that when $\alpha\geq 0.5$ no transaction is safe. With the results we provide, for example for Bitcoin which follows $k$-confirmation rule with $k=6$, even if $\gamma=0$, every transaction is at risk of double-spend for $\alpha>0.409$. This, however, does not mean, that the transaction is going to be replaced with a conflicting one without fail, instead, it means, that trying the double-spending attack with no-trailing comes at no-cost to the adversary. Moreover, the attacks have three possible outcomes from the perspective of the transaction:
  1. 1.
    (Service) The original transaction is confirmed and stays in the longest chain in the long run.
  2. 2.
    (Move-Funds) The original transaction is mined into a block, however, it is replaced by a conflicting transaction before being confirmed.
  3. 3.
    (Double-Spend) The original transaction is confirmed and later replaced by a conflicting transaction.
We note that if the conflicting transaction is a transaction, where adversary moves its money between sybil accounts, move-funds event can be associated with a second chance that adversary gets at double-spending. Moreover, for a blockchain with finite block size, it decreases the effective throughput. For each $\alpha$ and $\gamma$, we provide the success probability of each event above, which essentially shows how great that risk is. For example, the double-spending risk (i.e., the double-spending probability) for $k=6$, $\alpha=0.41$ and $\gamma=0$ is $0.092$. For $k=6$ and $\gamma=0$, as $\alpha$ approaches $0.5$, the double-spending risk approaches $0.21$. On the other hand, for $k=6$, $\alpha=0.41$ and $\gamma=1$, the double spending risk is $0.002$, whereas service probability is $0.59$. Finally, we note that the scheme we provide assumes no pre-mining, hence, the security threats could be much worse under settings with pre-mining.
We also provide a modified version of $L$-stubborn mining attack, which we call $S$-stealth mining, that is more similar to the classical private attack, in the sense that it does not expose itself by matching blocks and increases the success probability of double-spending attack with no-trailing at the cost of revenue and move-funds event. We analyze the properties of $S$-stealth mining as we do for the $L$-stubborn mining attack.
For both $L$-stubborn mining and $S$-stealth mining attacks, we provide a new metric for revenue calculation, where the extra revenue from the double-spending attack is taken into consideration in addition to the usual block rewards. Further, we also provide minimum transaction value for double-spending attacks that makes sure that the overall average revenue is larger than the honest mining protocol, when the usual revenue is less than the honest mining protocol.
In our analysis, we resort to the Bertrand’s ballot problem[24], Bailey’s number[25] and Catalan numbers[26], which have appeared explicitly in the blockchain literature in [27] where Chen et. al. calculate the probability that a chain wins a fork race given different mining probabilities depending on which fork is leading. Although Chen et. al. correctly argue that celebrated blockchain problems such as Nakamoto’s security problem [1] and selfish mining phenomenon [4] can be cast as modified ballot problems, they only provide the probabilities of each fork being ahead and their asymptotic values. Catalan numbers further appeared in [28, 29] where Grunspan and Marco make use of them to calculate the revenue of the stubborn mining [7] in closed form.
Our analysis considers average reward ratio MDP of the [6, 10] for no-trail stubbornness and we provide a novel algorithm to solve the problem, which outputs the optimal values in finite steps. Our approach can be a basis to solve the general MDP problem, i.e., with trail stubbornness, directly, which we leave for future research.
We summarize the contributions of our paper as follows:
  • Given pool size $\alpha$ and pool’s network influence $\gamma$, we find the optimal level of no-trail stubbornness $L$ and provide the associated revenue.
  • We connect the attack to double-spending, as $L>k$ implies double-spending. Thus, given $\alpha$ and $\gamma$, we provide values of $k$ where $k$-confirmation rule is at risk of double-spending (which comes at no-cost to the adversary).
  • We provide the double-spending probability, i.e., the risk, for each attack cycle.
  • We provide a modified version of the attack that works by stealth in order to prevent exposure and increase the double-spending probability at the cost of mining revenue.
  • We provide the minimum double-spend value needed for an attack to be profitable in the regimes where the scheme is less profitable than honest mining.
Different from [8, 9], we provide a simple formula to calculate revenues, double-spend risks and optimal attacks. As we do not deal with MDPs, the numerical results in Section 6 are extremely easy-to-analyze and provide a great insight at first sight. Moreover, the strategy we provide is easy-to-follow compared to complicated prescriptions of MDPs.
2 Preliminaries
2.1 Catalan Numbers and Dyck Words
Let $S$ be a string consisting of $X$’s and $Y$’s. If no prefix of $S$ has more $Y$’s than $X$’s, then we call $S$ a Pre-Dyck word. If a Pre-Dyck word has equal number of $X$’s and $Y$’s, it is called a Dyck word. Any Pre-Dyck word can be converted to a Dyck word by appending enough number of $Y$’s. The total number of Pre-Dyck words that contain $n$ $X$’s and $m$ $Y$’s is denoted as $P[n,m]$. The number of Dyck words of length $2n$, i.e., $P[n,n]$, is called Catalan number $C_{n}$ [26], and satisfies
$C_{n}=\frac{1}{n+1}{2n\choose n}.$ (1)
Catalan numbers also satisfy the following recursion
$C_{n}=\frac{2(2n-1)}{n+1}C_{n-1}$ (2)
with $C_{0}=1$. The generating function of the Catalan numbers is
$C(x)=\sum_{n\geq 0}C_{n}x^{n}=\frac{2}{1+\sqrt{1-4x}}.$ (3)
Every Pre-Dyck word starts with $X$ and every Dyck word ends with $Y$, i.e., $P[n,n]=P[n,n-1]=C_{n}$. A Pre-Dyck word can end with $Y$ if $m>0$ and with $X$ if $n>m$. Thus, Pre-Dyck words satisfy the recursion
$P[n,m]=P[n,m-1]+P[n-1,m].$ (4)
Clearly $P[n,m]=0$ if $m>n$ and it is easy to verify that $P[n,0]=1$ and $P[n,1]=n$. Thus, (4) gives a straightforward formula for finding $P[n,m]$ in $\mathcal{O}(nm)$. Every Dyck word with length $2n$ has a suffix that starts with $1$ $X$ and $r$ $Y$’s, which implies
$C_{n}=\sum_{m=0}^{n-1}P[n-1,m].$ (5)
2.2 Bailey’s Number and Bertrand’s Ballot Problem
Although it is easy to find a formula for $P[n,m]$ by converting (4) to frequency domain using generating functions, one could draw an analogy to the well-known Bertrand’s ballot problem [24], which asks the probability that a candidate A, who received $n$ votes, was strictly ahead of a candidate B, who received $m$ votes, throughout an election. By slightly modifying Andre’s solution to the ballot problem [30], which can be used to count the number of paths from $(a,b)$ to $(c,d)$ above and below $y=x$, to allow ties, we find
$P[n,m]=\frac{n-m+1}{n+m+1}{n+m+1\choose n+1}.$ (6)
Note that, if a broad range of $P[n,m]$ is needed, it is better to start from $P[1,0]=1$ and fill the matrix $P$ via (4).
The elements of $P$ matrix appear in [25] where Bailey investigates the sequences that contain $n$ $1$’s and $m$ $-1$’s such that the running sum is always non-negative. Bailey proved that the number of such sequences is equal to $P[n,m]$ provided in (6). Further, the diagoanal and antidiagonal (including all the below diagonal and antidiagonal) elements of matrix $P$ overlap with the columns of $A^{(1,2)}$ defined in [31] as Catalan-like numbers, which is not surprising since the recursion equations are similar.
2.3 Wald’s Identity and Doob’s Optional Stopping Time
Consider a Pre-Dyck word with $L$ $X$’s and $m<L$ $Y$’s. Append a random sequence of $X$’s (w.p. $\alpha$) and $Y$’s (w.p. $1-\alpha$) until the first time the number of $Y$’s becomes 1 less than the number of $X$’s. As this is a stopping time, by Wald’s identity, the expected length of such a sequence is $\frac{L-m-1}{1-2\alpha}$. As $L-m-1$ $Y$’s are required to reduce the gap to $1$, the rest of the characters in the string are split evenly between $X$ and $Y$. Thus, the final sequence contains $L+\frac{(L-m-1)\alpha}{1-2\alpha}$ $X$’s in expectation.
3 System Model
The setting we consider in this paper for the blockchain model is the same as the model considered in the seminal work on selfish mining by Eyal and Sirer[4]. We assume that there are three groups of miners. The first group can be considered as a single-entity which constitutes $\alpha=1-\beta$ fraction of the mining power in the system and acts adversarially, i.e., does not follow the longest-chain protocol and can decide to mine privately without sharing the blocks it has mined. The second and third groups together comprise the honest miners who follow the longest-chain protocol, i.e., mine on top of the longest-chain they have seen and release the newly mined blocks immediately. The former of the two constitutes $\gamma\beta$ fraction of the mining power is influenced by the adversarial group and accepts the adversarial chain in case of a tie, whereas the latter constitutes $(1-\gamma)\beta$ fraction of the mining power and accepts the honest chain in case of a tie. The reason for such a partition and in which real-world scenarios this can arise is explained in [4].
We only consider the blocks that contain a valid PoW (Proof-of-Work) and assume that the inter-arrival times between blocks follow an exponential distribution. Further, block propagation delays are assumed to be negligible, hence no blocks are mined while a new block is being propagated through the network. As a result, each mined block belongs to one of the three groups with a probability proportional to their mining power. Moreover, the blockchain protocol updates the puzzle difficulty threshold so that the average inter-arrival time between blocks is constant. Each block rewards its miner a predetermined amount of coins, i.e., coinbase, which we assume is constant for the foreseeable future. If each miner acts honestly, the coinbase reward they get is proportional to their mining power in the network in the limit. However, adversarial strategies which deviate from the longest-chain protocol result in rewards disproportional to the mining powers. We refer to the adversarial strategies that aim to increase their proportion of the coinbase rewards as selfish mining strategies. Although these selfish mining strategies initially decrease the rewards, they can be seen as an attack on the block interval adjustment [32], which in turn increases the rewards in the long run.
A transaction $tx$ in this system is considered to be confirmed according to a $k$-deep confirmation rule if a block that contains $tx$ has $k-1$ blocks mined on top of it and is part of the longest chain available to the honest miners. Double-spending refers to the event that $tx$ is confirmed according to the $k$-deep confirmation rule and a conflicting transaction $tx^{\prime}$ replaces $tx$ in the longest chain and becomes $k$-deep.
3.1 $L$-Stubborn Mining
The selfish mining strategies are extensively analyzed in the literature with notable works such as [4, 7, 6]. In this work, we consider a sub-group of strategies from all possible states and action space, which we call $L$-stubborn [^1] In this paper, $L$-stubborn mining is different than the lead-stubbornness (L-stubbornness) defined in [7]. There, “L” stands for the word “lead.” Here, $L$ refers to an integer which is the length of the adversarial chain with respect to a common ancestor block between the honest and the adversarial chains. mining. To describe the strategy, let us assume that all participants start with the same chain, which we call the offset chain. Both the adversary and the honest miners mine on top of the offset chain. If the honest miners are the first to mine, then the adversary accepts the new block and redefines the offset chain with the new block on its tip. For the sake of simplicity, we denote the height of the offset chain as zero. If the adversary is the first to mine a block on top of the offset chain, i.e., at height $1$, it keeps the block private and we denote the highest block height of the adversarial chain as $A$ and call the adversarial chain as $A$-chain. At any time later, if the honest miners mine a block on that height, adversary publishes its block on that height immediately if $A<L$, which we call matching. In accordance with the network model defined earlier, $\gamma$ proportion of the honest miners accept and mine on top of the adversarial block instead of the honest block. We denote the highest honest block height as $H$ and the chain starting from the offset and ending in the honest block at height $H$ as $H$-chain. The private stubborn mining strategy of the adversary continues on for all heights until $A=L$, i.e., whenever $L>A\geq H$ adversary immediately matches the block at the height of the newly mined honest block. If honest miners manage to mine a block on a height before the adversary before $A=L$, i.e., $L\geq H=A+1$, the adversary accepts the $H$-chain and redefines the offset chain. On the other hand if $A\geq L$, then the adversary keeps mining privately until $A=H+1$, at which point, it releases the $A$-chain and redefines the offset chain. An attack cycle is a process that starts when an offset chain is defined, and ends when a new offset chain is defined.
Note that, for $L=1$ the strategy reduces to honest mining, whereas for $L=2$ the strategy reduces to the celebrated selfish mining attack of [4]. On the other hand, for $L=\infty$ the strategy becomes the equal fork stubborn mining strategy described in [7] and analyzed in closed-form with Catalan numbers in [28]. In this paper, we are interested in giving a method for calculating the revenues for any given $L$ as well as deriving a method to find the best possible $L$ given $\alpha$ and $\gamma$. The resulting strategy does not allow any trail stubbornness discussed in [7], i.e., the adversary accepts the honest chain whenever it falls behind. Hence, we can only claim the optimality of the strategy under action spaces with independent attack cycles that allow no-trail stubbornness. We argue that the parameter $L$ represents the level of stubbornness in selfish mining and our analysis provides the optimal level of stubbornness as well as the maximum stubbornness such that the attack is still more profitable than honest mining. Thus, we view $L$ (and $S$, which will be defined later) as a parameter of stubbornness.
In [6], the authors use an undiscounted average reward MDP to provide $\epsilon$-optimal policies. $L$-stubborn mining allows a sub-group of strategies from all possible states and action space and we give an easy-to-evaluate formula to find optimal $L$. The authors of [6] further argue about combining double-spending attack with selfish mining attack, however, they do not provide any concrete method or analysis. Later, in [9], the same authors provide an MDP that accounts for rewards of double-spending attack, however, the MDP ignores rewards from adversarial blocks that do not result in double-spending. Thus, the optimal policies resulting from the MDP therein are not optimal in the sense that they provide maximum mining rewards with free double-spending. Further, the results and strategies provided in [9, 6, 8] are complicated with no easy-to-follow structure. $L$-stubborn mining that we consider in this paper is a policy which results in higher mining rewards than honest mining (as well as optimal ones) and has strong implications on double-spending strategies with a clear and easy-to-follow analysis under $k$-confirmation rule, which we explain next.
3.2 $k$-Confirmation Rule
First, we rigorously define the events related to a double-spending attack.
Definition 1
Consider a transaction $tx$ where the adversary pays a merchant to get a service and a conflicting transaction $tx^{\prime}$ where the adversary simply changes the address of the same funds, which it can use again later. We consider three disjoint events regarding $tx$ and $tx^{\prime}$:
  1. 1.
    Service If $tx$ is confirmed and makes it into the offset chain, we call the event Service.
  2. 2.
    Move-Funds If $tx$ is never confirmed, we call the event Move-Funds.
  3. 3.
    Double-Spending If $tx$ is confirmed but does not make it into the offset chain, we call the event Double-Spending.
Notice the relation between $L$-stubborn mining and $k$-confirmation rule. If $(k+1)$-stubborn mining attack is more profitable than honest mining, adversary can try to double-spend at no-cost [^2] As observed by [6], $tx$ is a real transaction and its confirmation results in a service, therefore, this should not be seen as a loss even from the perspective of the adversary. while implementing the $(k+1)$-stubborn mining strategy. All it needs to do is to include a $tx^{\prime}$ in the first height above the offset chain, whereas honest miners will include $tx$ on the first block above the offset chain. Later on, when $A$-chain is released, if $H$-chain contains $tx$, it will be replaced by $tx^{\prime}$, allowing the adversary to double-spend.
Before the offset event happens, if $A>k$, this implies that at some point $H=k$ and adversary will release $A$-chain later on. Depending on whether $H$-chain contains $tx^{\prime}$ or $tx$, either Move-Funds or Double-Spending happens. On the other hand, before the offset event happens, if $A<H\leq k$, depending on whether $H$-chain contains $tx^{\prime}$ or $tx$, either Move-Funds or Service happens. If an attack cycle ends with Move-Funds event, later, adversary can try double-spending again using the same funds. This also decreases effective throughput in finite block size systems. One could also rescale the probabilities of Double-Spending and Service by excluding Move-Funds event from the perspective of the actual funds associated with $tx$ and $tx^{\prime}$ since eventually the fund is going to result in Service or Double-Spending. We omit the rescaling in our calculations and numerical results, as rescaling is straightforward to do and keeping Move-Funds event gives an understanding of throughput decrease to the reader.
We also note that the time $tx$ is released in the network is another important issue while considering double-spending attacks. In general, it can be shown that a double-spending attack can be successful with probability $1$ if adversary is allowed to pick the time $tx$ is released, since it can build an arbitrarily high pre-mining gain [9]. The pre-mining gain refers to the difference between the private chain of the adversary and the public honest chain, in general [9]. The approach we take here assumes that the $tx$ is released at the start of each attack cycle, hence the pre-mining gain is 00. Thus, the results we provide are conservative and the adversary could get even more profits from double-spending with more clever approaches. Indeed, as we will see later in the numerical results section, the service probability is always above $\beta$, since the first block mined by the honest nodes contains $tx$ and with probability $\beta$ it is mined before any adversarial blocks and immediately becomes part of the offset chain.
3.3 $S$-Stealth Mining
Notice the difference between Nakamoto private attack and $(k+1)$-stubborn mining. The first major difference is $(k+1)$-stubborn mining gives up the attack when $A$-chain falls behind $H$-chain, i.e., no-trailing behind is allowed. Another major difference is, during a cycle, $(k+1)$-stubborn mining matches each honest block with an adversarial one which can expose the attack as two chains with conflicting transactions $tx$ and $tx^{\prime}$ are competing openly in the network, arousing the merchant’s suspicion. The matching also decreases Double-Spending probability while increasing Move-Funds probability, which we show later.
We address some of these issues by modifying $L$-stubborn mining in a way that only allows matching if both double-spending and the risk of losing mining revenue are imminent, i.e., $H=A=k=L-1$. We call this attack $S$-stealth mining, which we investigate in Section 5. We note that $S$-stealth mining does not allow any trailing behind either, hence, there is still a difference between a private attack and $S$-stealth mining.
4 Revenue & Double-Spending of $L$-Stubborn Mining
Instead of building a complex Markov chain and going through the tedious analysis of steady-state distribution, we can analyze the distinct possible paths an attack cycle can take. An attack cycle starts when an offset chain is defined, and ends when a new offset chain is defined. If during an attack cycle, the length of the $A$-chain reaches $L$, i.e., $A=L$, then we call the attack successful. On the other hand, if the length of the $H$-chain surpasses the length of the $A$-chain before $A$ reaches $L$, we call the attack cycle unsuccessful.
In Figure 1, we display the sample paths of $4$-stubborn mining, i.e., $L=4$. The $y$-axis and $x$-axis represent $A$ and $H$, i.e., the length of $A$-chain and $H$-chain, respectively. An attack starts from point $(0,0)$ and arrivals of adversarial (honest resp.) blocks are represented with light-green (orange resp.) arrows. A red arrow shows that the arrival of honest block has ended the cycle. A blue (red resp.) cross represents the endpoint of a successful (unsuccessful resp.) attack cycle. A dark-green arrow represents the arrival of an adversarial block that guarantees a successful attack cycle. The dark-green crosses represent the first time cycles are guaranteed to be successful and not ended yet. The cycles that pass through the dark-green crosses eventually end with an honest arrival in blue cross points $(x,x-1)$ where $x>4$, i.e., when the difference of chain lengths reduces to $1$.
Note that there is a third dimension of Figure 1, which tracks the number of adversarial blocks in the prefix of $H$-chain since every arrival of honest block has two choices of chains to extend. Further, the dynamics of this dimension is Markovian since each honest arrival either increases the previous value to $H-1$ (w.p. $\gamma$) or stays the same (w.p. $1-\gamma$) and does not depend on anything else. This Markovian structure implies that we can separate the third dimension and treat it independently which simplifies our analysis, hence, we do not plot it.
Figure 1: Sample paths of $4$-stubborn mining.
Lemma 1
Under the $L$-stubborn mining strategy, the probability that $A$-chain reaches length $L$ during an attack cycle while the largest honest block is at height $m<L$ is
$P_{s}(L,m)=\frac{L-m}{L+m}{L+m\choose L}\alpha^{L}\beta^{m}.$ (7)
Proof:  We know that the winner of the cycle is determined once the last adversarial block arrives, hence we distinguish it from the previous $L-1$ blocks to prevent double-counting. The number of distinct paths with $H=m<L$ blocks and $A=L$ blocks, where the last block is adversarial, is equal to $P[L-1,m]$; see Figure 1 for an example where $L=4$. $\blacksquare$
Lemma 2
The probability that an attack cycle is unsuccessful and ends with the largest honest block at height $n+1$ which has $i\leq n$ adversarial blocks in its prefix (after the offset chain) is
$P_{u}(n,i)=\frac{1}{n+1}{2n\choose n}\alpha^{n}\beta^{n+1}(1-\gamma)^{n-i}\gamma^{\mathbbm{1}_{i}}.$ (8)
Proof:  The end of the unsuccessful attack cycle implies that the last block that arrives has to be honest and the adversary was never behind up until that the last block. As a result, the arrival sequence of honest and adversarial blocks creates a Dyck string with length $n<L$. There are $C_{n}$ number of such strings and together with the last honest block each has probability $\alpha^{n}\beta^{n+1}$. Hence, an attack cycle ends unsuccessfully with the honest block at height $n+1$ with probability
$P_{u}(n)=\frac{1}{n+1}{2n\choose n}\alpha^{n}\beta^{n+1}.$ (9)
Further, each time an honest block arrives, it is either mined by the honest miners who are under the influence of adversarial view w.p. $\gamma$, or it is mined on top of the last honest block. As a result, no adversarial block is contained in the prefix of the honest block on height $n+1$ w.p. $(1-\gamma)^{n}$ or $i>0$ adversarial blocks are contained w.p. $(1-\gamma)^{n-i}\gamma$. Note that, $\mathbbm{1}_{i}=1$ if $i\neq 0$ and zero otherwise. $\blacksquare$
Next, we put Lemma 1 and Lemma 2 together to obtain the revenue of selfish mining.
Theorem 1
(Revenue) The revenue of $L$-stubborn mining is given as
$\rho_{L}=\frac{L_{s}+L_{u,a}}{L_{s}+L_{u}},$ (10)
where
$L_{s}$ $=\sum_{m=0}^{L-1}P_{s}(L,m)\left(L+\frac{(L-m-1)\alpha}{1-2\alpha}\right),$ (11)
$L_{u,a}$ $=\sum_{n=0}^{L-1}\sum_{i=0}^{n}P_{u}(n,i)i,$ (12)
$L_{u}$ $=\sum_{n=0}^{L-1}P_{u}(n)(n+1).$ (13)
Proof:  In Lemma 1 and Lemma 2, we provide the probability for each possible outcome of an attack cycle. If an attack cycle is successful, we have an $A$-chain at length $L$ and the largest honest block is at height $m<L$. Clearly, if $L-m>1$, the adversary will continue the cycle until the difference goes down to $1$. Using Wald’s identity provided in Section 2.3, the expected revenue in this case is $L+\frac{(L-m-1)\alpha}{1-2\alpha}$. This explains $L_{s}$ in (10). On the other hand, if the attack cycle is unsuccessful, $i$ adversarial blocks still make it into the offset chain w.p. $P_{u}(n,i)$ and the total number of blocks that make it into the offset chain is $n+1$. As a result, these two cases are considered in the numerator and denominator of (10) as $L_{u,a}$ and $L_{u}$, respectively. $\blacksquare$
Note that, a successful attack cycle does not immediately imply a successful double-spending. It only implies that the $A$-chain will make it into the offset chain eventually. Similarly, an unsuccessful attack cycle does not immediately imply that $tx$ will be confirmed.
Next, we provide the probability of each event regarding $tx$ and $tx^{\prime}$ for $(k+1)$-stubborn mining with $k$-confirmation rule.
Theorem 2
[…]
$P(\textbf{Service})$ $=\sum_{m=0}^{k}P_{u}(m)(1-\gamma)^{m},$ (15)
[…]
Proof:  For Double-Spending to happen, one possibility is, we will observe $A>k$ and $H\leq k$ at some point during the cycle. The first time this happens during the cycle is the first time $A=k+1$ and $H=m$ which happens w.p. $P_{s}(k+1,m)$. Until this point, adversary matches every one of the $m$ honest blocks. From this point onwards, adversary does not need to match the honest blocks as the strategy will eventually publish the longer $A$-chain. Hence, $tx$ stays in the $H$-chain if none of the $m$ matching adversarial blocks make it into the prefix of the the honest block at height $H$, which happens with probability $(1-\gamma)^{m-\mathbbm{1}_{m=k}}$. Note that, if $H=k$, then it does not matter whether prefix chain is after the $k$th block or not, hence $\mathbbm{1}_{m=k}$ at the exponent. Another possibility for Double-Spending is, if $H$-chain catches $A$-chain with both having length $k$ and the next block is honest which has probability $P_{u}(k)$. In this situation, if $H$-chain switches its prefix on the last honest block from $tx$ to $tx^{\prime}$, which happens w.p. $(1-\gamma)^{k-1}\gamma$, then $tx$ gets discarded despite being confirmed earlier. For Service to happen, we will observe $L\geq H=A+1=m+1$, which ends the cycle with $H$-chain making it into the offset chain which happens w.p. $P_{u}(m)$. Further, $tx$ has to be in the prefix of the honest block at height $m+1$, which explains the term $(1-\gamma)^{m}$. $\blacksquare$
4.1 Double-Spending Danger is Imminent Well Before $\alpha=0.5$
We define
$L^{*}$ $=\arg\max_{L}\rho_{L},$ (17)
$\bar{L}$ $=\sup\{L:\rho_{L}\geq\alpha\},$ (18)
where $\bar{L}\geq L^{*}$ due to the quasiconcavity proven in Theorem 3. Essentially, any $k$-confirmation rule with $k<\bar{L}$ is compromised by $(k+1)$-stubborn mining since adversary gains more revenues than honest mining, i.e., $\rho_{k+1}\geq\alpha$, and at the same time it can double-spend at no-cost. In other words, for a given $\alpha$ and $\gamma$, $k$-confirmation rule should be at least $k\geq\bar{L}$ to make sure that double-spending attack has revenue loss. Note that, if $\rho_{\infty}\geq\alpha$, no $k$-confirmation is safe as double-spending has zero-cost. From this perspective, the safety of a blockchain system fails well before adversarial proportion reaches $\alpha$. We investigate this phenomenon further in the numerical results section by finding $L^{*}$ and $\bar{L}$ for each $\alpha$ and $\gamma$.
To find the optimal $L$, one could evaluate $\rho_{L}$ in (10) and find the $L$ that provides the highest revenue. As the calculations involve binomial coefficients and (10) involves fractions, it is not straightforward to find $L^{*}$. Instead, we provide a much simpler method to obtain the optimal $L$. To do so, we resort to the linearization method of the MDP problem [6] and use it in a narrow-sense. Note that, every time a block is mined, the adversary faces a decision as to whether it should publish $A$-chain and end the cycle, or continue. Clearly, if $A-H>1$, the decision should be to continue the attack cycle as the adversary can gain additional advantage at no-cost in these situations. On the other hand, if $A=H+1$, the risk is imminent as the adversary loses the advantage of overwriting the $H$-chain. As a result, we can narrow down the decision problem to a specific case where $A=H+1$ and analyze the situation therein via the linearization method.
To that end, we define
$w_{\rho}(x,y)=(1-\rho)x-\rho y,$ (19)
as in [6], and $u$ and $v_{a}$ as
$u=(1-\gamma),\qquad v_{a}=\frac{1}{u}\left(1-\gamma\frac{1-\rho_{a}}{1-2\alpha}\right).$ (20)
Next, using the definitions above and intermediary results in Appendix A, we provide the following quasiconcavity result related to the $L$-stubborn mining.
Theorem 3
(Quasiconcavity) $\rho_{L}\geq\rho_{L+1}$ implies $\rho_{a}\geq\rho_{a+1}$ for any $a\geq L$. Similarly, $\rho_{M}\geq\rho_{M-1}$ implies $\rho_{a}\geq\rho_{a-1}$ for any $a\leq M$.
Theorem 3 essentially says that $\rho_{L}$ is quasiconcave for integer $L$, which implies if $L^{*}>k+1$, $(k+1)$-stubborn mining is still more profitable than honest mining. We relegate the proof of Theorem 3 to the Appendix A as it requires some intermediary definitions and results. These results also shed light on a simple algorithm that can be used to find $L^{*}$.
Next, we provide Algorithm 1 that finds the optimal $L^{*}$ of $L$-stubborn mining. For that, let $\rho_{1}$, $\rho_{2}$ and $\rho_{\infty}$ be defined as follows
$\rho_{1}$ $=\alpha,$ (21)
$\rho_{2}$ $=\frac{\alpha\beta^{2}(4\alpha+\gamma(1-2\alpha))-\alpha^{3}}{1-\alpha(1+(2-\alpha)\alpha)},$ (22)
$\rho_{\infty}$ $=\frac{\alpha}{\beta}-\frac{(1-2\alpha)(1-\gamma)}{\beta\gamma}(1-\beta C((1-\gamma)\alpha\beta)),$ (23)
which are the revenues of honest mining, selfish mining [4], and equal-fork stubborn mining [7, 28], respectively. Note that each of these values can be obtained by (10), where for $\rho_{\infty}$, $L_{s}$ can be set to zero as the $H$-chain will catch the $A$-chain eventually, i.e., no attack ends successfully, but revenue comes from matching. In Appendix A, we use the intermediary results to show that Algorithm 1 below outputs $L^{*}$.
Algorithm 1 Optimal $L^{*}$ algorithm
1:  $i\leftarrow 1$
2:  $L^{(1)}\leftarrow\arg\max(\rho_{1},\rho_{2},\rho_{\infty})$
3:  if $v_{L^{(1)}}\leq 0$ then
4:     return  $\infty$
5:  end if
6:  while $L^{(1)}>1$ do
7:     $L^{(i+1)}\leftarrow\lceil\frac{\log v_{L^{(i)}}}{\log u}\rceil$
8:     if $L^{(i+1)}=L^{(i)}$ then
9:        return  $L^{(i)}$
10:     else
11:        $i\leftarrow i+1$
12:     end if
13:  end while
14:  return  $L^{(i)}$
4.2 Combined Revenue
In this section, we provide a formula that incorporates the profit of double-spending into the revenue the adversary makes from $(k+1)$-stubborn mining. To do so, we make the assumption of [9], i.e., we assume that the profit of every successful double-spend is equal to $\mathcal{R}$ units of block rewards. Further, as in [9], we consider the replacement of every confirmed block as different instances of double-spend events, i.e., a cycle where $l$ honest blocks are confirmed and later replaced, result in extra $l\cdot\mathcal{R}$ revenues in block reward units. Notice that, in this paper, we treat the revenue as the relative adversarial block rewards in the system. Thus, we also treat the double-spending revenue as extra block rewards the adversary can get in addition to the actual block rewards. Other considerations are left as simple extensions which the reader can easily calculate given the approaches we provide here.
Theorem 4
(Combined Revenue) The combined revenue of $(k+1)$-stubborn mining and the associated double-spending is given as
$\rho^{\prime}_{(k+1)}(\mathcal{R})=\frac{(k+1)^{\prime}_{s}+(k+1)^{\prime}_{u,a}}{(k+1)_{s}+(k+1)_{u}},$ (24)
where $(k+1)_{s}$, $(k+1)_{u,a}$ and $(k+1)_{u}$ are defined as in Theorem 1, where we pick $L=(k+1)$ and
$(k+1)^{\prime}_{s}$ $=(k+1)_{s}+\mathcal{R}\cdot\sum_{m=0}^{k}P_{s}(k+1,m)(1-\gamma)^{m-\mathbbm{1}_{m=k}}\left(1+\frac{(k-m)\alpha}{1-2\alpha}\right),$ (25)
$(k+1)^{\prime}_{u,a}$ $=(k+1)_{u,a}+\mathcal{R}\cdot P_{u}(k)(1-\gamma)^{k-1}\gamma.$ (26)
Proof:  Note that the first term in (14) provides the sum of successful cycles with double-spending events. Each term in that sum corresponds to the first time $A=k+1$ and $H=m$, and none of the matched adversarial blocks make it into the prefix of the honest block as explained in the proof of Theorem 2. Given Wald’s identity and the fact that a cycle will continue until the difference goes down to $1$, all but last $k$ adversarial blocks will result in double-spending events. Thus, total extra block rewards adversary gets in a successful attack cycle in addition to the $(k+1)_{s}$ is provided as in (25), where the extra term is $\mathcal{R}$ times the expected number of honest blocks that are confirmed and later replaced in an attack cycle. An unsuccessful cycle can only result in double-spending if the event corresponding to the second term in (14) happens, where only one honest block is confirmed and replaced later, which explains the extra term in (26). $\blacksquare$
We define
$R^{*}_{k}$ $=\inf\{\mathcal{R}:\rho^{\prime}_{k+1}(\mathcal{R})\geq\alpha\},$ (27)
which, essentially, gives the minimum double-spending reward value required for the scheme to be profitable on average. We provide $R^{*}_{k}$ for each $\alpha$ and $\gamma$ value for Bitcoin, where $k=6$, in numerical results later.
Next, we provide another approach that incorporates the profit coming from double-spending when considering the $(k+1)$-stubborn mining attack provided here. As mentioned in [3], the value of service provided might be below the fee paid. In this case, we assume that each successful cycle results in only single double-spending event, i.e., only the first block in the $H$-chain is attacked which contains $tx$. We further assume that the value of service provided by the merchant equals $V$ and adversary pays $F$ fee for the service with $F\geq V$, where both values are in block reward units. In this case, on average, the $(k+1)$-stubborn mining attack is more profitable than honest mining if
[…]
$\geq\alpha\cdot\left((k+1)_{s}+(k+1)_{u}\right)+(F-V)\cdot P(\textbf{Service}).$ (28)
The reasoning for this is as follows: Out of every $\left((k+1)_{s}+(k+1)_{u}\right)$ blocks that make it into the prefix chain, if the adversary followed the honest protocol, $\alpha$ proportion would be its profit. However, with the $(k+1)$-stubborn mining protocol, $(k+1)_{s}+(k+1)_{u,a}$ blocks make it into the prefix chain. Further, the double-spending results in $V$ extra rewards for the adversary. However, if only Service happens, $F-V$ amount of rewards is lost.
5 $S$-Stealth Mining
[…]
In $S$-stealth mining, the strategy of keeping track of the length of $A$-chain is still the same, however, matching only happens if $A=H=S-1$. In other words, whenever the honest miners mine a new block on a height $H<S-1$, the adversary does not match the honest block. On the other hand, if $A=S-1$ and $H=S-2$ and the next block is honest on height $S-1$, the adversary matches this block.
Theorem 5
(Revenue) The revenue of $S$-stealth mining is given as
$\sigma_{S}=\frac{S_{s}+S_{u,s}}{S_{s}+S_{u}},$ (29)
where
$S_{s}$ $=\sum_{m=0}^{S-1}P_{s}(S,m)\left(S+\frac{(S-m-1)\alpha}{1-2\alpha}\right),$ (30)
$S_{u,s}$ $=P_{u}(S-1)\gamma(S-1),$ (31)
$S_{u}$ $=\sum_{n=0}^{S-1}P_{u}(n)(n+1).$ (32)
Proof:  For a successful attack, we follow the same arguments provided in the proof of Theorem 1. When an attack is unsuccessful, adversarial blocks make it into the offset chain only if the $H$-chain has reached length $S-1$, and the next honest block on height $S$ switches its prefix (w.p. $\gamma$) to the $A$-chain which has length $S-1$. $\blacksquare$
We note that $\sigma_{1}=\rho_{1}=\alpha$ and $\sigma_{2}=\rho_{2}$, i.e., $2$-stealth mining is the same algorithm as selfish mining of [4]. Further, $\sigma_{\infty}=0$ as there is no matching when $S=\infty$ and eventually $H$-chain will overtake in every cycle.
Theorem 6
[…]
Proof:  Contrary to $(k+1)$-stubborn mining, a successful cycle of $(k+1)$-stealth mining always results in double-spending since the $H$-chain always keeps $tx$ in its prefix and the $A$-chain undoes it eventually. On the other hand, if an attack ends unsuccessfully with $H=(k+1)$, the honest chain might switch from $tx$ to $tx^{\prime}$ right at the end when the honest block at height $k+1$ is mined. This would imply that the first $tx$ is confirmed when the $H$-chain reaches length $k$ and the next honest block switches prefix and $tx^{\prime}$ is confirmed, resulting in a double-spend. $\blacksquare$
Notice that $(k+1)$-stealth mining is essentially the private attack of Nakamoto [1] where the adversary gives up pursuing the attack if it falls behind before reaching length $k$. If it reaches length $k$ but the honest chain catches it at length $k$, the network split $\gamma$ could still help the adversary to double-spend. However, as matching only happens at length $k$, there is no Move-Funds event possible in $(k+1)$-stealth mining.
The formula we provide in Theorem 5 can be used to determine the revenue of $(k+1)$-stealth mining which is no different than the private attack with no-trail. If $\sigma_{k+1}\geq\alpha$, this would imply that trying private attack with no-trail comes at no-cost, i.e., mining revenues are still above or equal to $\alpha$. Hence, similar to the case of $L$-stubborn mining we define
$S^{*}$ $=\arg\max_{S}\sigma_{S},$ (35)
$\bar{S}$ $=\sup\{S:\sigma_{S}\geq\alpha\},$ (36)
and provide their associated values for given $\alpha$ and $\gamma$ in the numerical results section. For the sake of completeness, we provide here the quasiconcavity of $S$-stealth mining and an algorithm to find $S^{*}$. The proofs are provided in Appendix B.
Theorem 7
(Quasiconcavity) $\sigma_{S}\geq\sigma_{S+1}$ implies $\sigma_{a}\geq\sigma_{a+1}$ for any $a\geq S$. Similarly, $\sigma_{R}\geq\sigma_{R-1}$ implies $\sigma_{a}\geq\sigma_{a-1}$ for any $a\leq R$.
Next, we provide Algorithm 2 below, which returns $S^{*}$, where
$f(\sigma)=\sup\left\{x\in\mathbb{R}:1-\frac{\gamma(1-2\alpha)}{\alpha(1-\alpha)}\frac{x^{2}-1}{2(2x-1)}-(1-2\alpha)(x+1-\gamma x)\geq\sigma\right\},$ (37)
and note that $\sigma_{1}=\alpha$ and $\sigma_{2}=\rho_{2}$, which is provided in (22).
Algorithm 2 Optimal $S^{*}$ algorithm
1:  $i\leftarrow 1$
2:  $S^{(1)}\leftarrow\arg\max(\sigma_{1},\sigma_{2})$
3:  while $S^{(1)}>1$ do
4:     $S^{(i+1)}\leftarrow\lceil f(\sigma_{S(i)})\rceil$
5:     if $S^{(i+1)}=S^{(i)}$ then
6:        return  $S^{(i)}$
7:     else
8:        $i\leftarrow i+1$
9:     end if
10:  end while
11:  return  $S^{(i)}$
5.1 Combined Revenue
Here, we first provide the approach where each honest block that is confirmed but does not make it into the offset chain, results in $\mathcal{R}$ additional block rewards for the adversary. In this case, the revenue is provided in Theorem 8.
Theorem 8
(Combined Revenue) The combined revenue of $(k+1)$-stealth mining and the associated double-spending is given as
$\sigma^{\prime}_{(k+1)}(\mathcal{R})=\frac{(k+1){{}^{\prime\prime}}_{s}+(k+1){{}^{\prime\prime}}_{u,s}}{(k+1)_{s}+(k+1)_{u}},$ (38)
where $(k+1)_{s}$ and $(k+1)_{u}$ are defined as in Theorem 5, where we pick $S=(k+1)$ and
$(k+1){{}^{\prime\prime}}_{s}$ $=\sum_{m=0}^{k}P_{s}(k+1,m)\left(k+(\mathcal{R}+1)\left(1+\frac{(k-m)\alpha}{1-2\alpha}\right)\right),$ (39)
$(k+1){{}^{\prime\prime}}_{u,s}$ $=P_{u}(k)\gamma(\mathcal{R}+k).$ (40)
Proof:  Since there is no matching in $(k+1)$-stealth mining except when both $A=H=k$, all successful attack cycles end up with double-spending, where when the $A$-chain makes into the offset chain all adversarial blocks except the last $k$ result in double-spending which explains the $\mathcal{R}$ factor in (39). As explained in the proof of Theorem 6, an unsuccessful attack cycle can result in double-spending only if the prefix switch happens at the last honest mining which explains (40). $\blacksquare$
We define the minimum double-spend reward value needed for a profitable scheme as
$R^{*}_{k}$ $=\inf\{\mathcal{R}:\sigma^{\prime}_{k+1}(\mathcal{R})\geq\alpha\}.$ (41)
If we assume that each successful cycle results in only a single double-spending event with a service value $V$ and payment $F$ with $F\geq V$ both in block rewards units, then, on average, the $(k+1)$-stealth mining is more profitable than honest mining if
$(k+1)_{s}+(k+1)_{u,s}+V\geq\alpha\cdot\left((k+1)_{s}+(k+1)_{u}\right)+F\cdot P(\textbf{Service}),$ (42)
where all values of $(k+1)_{s}$, $(k+1)_{u,s}$, $(k+1)_{u}$, $P(\textbf{Service})$ are provided as in Theorem 5 and Theorem 6.
(a) $L^{*}$
(b) $S^{*}$
(c) $\frac{\rho_{L^{*}}}{\alpha}$
(d) $\frac{\sigma_{S^{*}}}{\alpha}$
(e) $\bar{L}$
(f) $\bar{S}$
Figure 2: $L$-stubborn mining and $S$-stealth mining.
(a) P(Double-Spending)
(b) P(Move-Funds)
(c) P(Service)
(d) $\frac{\rho_{k+1}}{\alpha}$
Figure 3: $(k+1)$-stubborn mining with $k=6$.
(a) P(Double-Spending)
(b) P(Service)
(c) $\frac{\sigma_{k+1}}{\alpha}$
Figure 4: $(k+1)$-stealth mining with $k=6$.
(a) Stubborn mining
(b) Stealth mining
Figure 5: $R^{*}_{6}$ for Bitcoin
6 Numerical Results
We start by evaluating $L^{*}$ for each value of $\alpha$ and $\gamma$ using Algorithm 1. The results are provided in Figure 2(a). Note that the boundary between $L^{*}=1$ and $L^{*}=2$, i.e., the honest mining and selfish mining, was already provided in [4]. Our results here show new boundaries by providing regions where waiting for the private branch to extend further is beneficial in terms of the revenue. Next, in Figure 2(c) for each $\alpha$ and $\gamma$, we provide the relative revenue of optimal $L^{*}$-stubborn mining attack compared to honest mining. For all values of $\gamma$ the relative revenue starts as $1$ from $\alpha=0$ and reaches $2$ at $\alpha=0.5$. However, as $\gamma$ increases, the revenue increases faster initially. Finally, in Figure 2(e), we provide $\bar{L}$ for each $\alpha$ and $\gamma$, which we believe has a significant meaning related to $(k+1)$-confirmation rule. As we proved that $\rho_{L}$ has a quasiconcave behavior, if $\bar{L}>k$, then $(k+1)$-stubborn mining is more profitable than honest mining, hence following the $(k+1)$-stubborn mining allows double-spending at no-cost. This in turn means, for each $\alpha$ and $\gamma$, $k$-confirmation rule has to be set according to $k\geq\bar{L}(\alpha,\gamma)$ to minimize the risk of double-spending that comes at no-cost to the attacker.
Similar to the results of $L$-stubborn mining, Figure 2 also contains related results for $S$-stealth mining. Observing the $S^{*}$ provided in Figure 2(b), it is clear that $S^{*}\leq L^{*}$ which is straightforward, but also $S^{*}\leq 4$ for $\alpha<0.5$ and $\gamma<1$. In other words, in stealth mining, adversary cannot act in a stubborn manner forever, whereas stubborn mining is the best strategy for selfish mining in the white region of Figure 2(a). In Figure 2(d), we provide the relative revenue of optimal $S^{*}$-stealth mining attack compared to honest mining. Although it starts as $1$ from $\alpha=0$ and reaches $2$ at $\alpha=0.5$ similar to the $L^{*}$-stubborn mining, the rise in revenue is not as steep initially. Further, as expected, the revenue difference between $L^{*}$-stubborn mining and $S^{*}$-stealth mining increases with $\gamma$, as matching is only present in the last step in stealth mining. Finally, Figure 2(f) shows the values of $\bar{S}$, where as $\sigma_{\infty}=0$, we have $\bar{S}<\infty$ for all values of $\alpha<0.5$ and $\gamma<1$.
[…]
For $(k+1)$-stubborn mining, the double-spending risk for $k=6$, $\alpha=0.41$ and $\gamma=0$ is $0.092$. For $k=6$ and $\gamma=0$, as $\alpha$ approaches $0.5$, the double-spending risk approaches $0.21$. On the other hand, for $k=6$, $\alpha=0.41$ and $\gamma=1$, the double-spending risk is $0.002$, whereas service probability is $0.59$. For $\gamma=0$, the same values apply to $(k+1)$-stealth mining as well. On the other hand, for $(k+1)$-stealth mining with $k=6$, $\alpha=0.41$ and $\gamma=1$, the double-spending risk is increased to $0.108$, whereas service probability increases to $0.892$. As mentioned earlier, the cost is a cut in relative revenue, which drops from $\frac{\rho_{7}}{\alpha}=1.639$ to $\frac{\sigma_{7}}{\alpha}=1.09$.
[…]
7 Conclusion and Extensions
In this paper, we provided a rigorous and explicit stochastic analysis for the combination of the two celebrated attacks: double-spending and selfish mining. To do so, we first combined stubborn and selfish mining attacks, i.e., constructed a strategy where the attacker acts stubborn until its private branch reaches a certain length $L$, and then switches to act selfish. We provided an analysis of the optimal stubbornness for each parameter regime as well as the maximum stubbornness that is still more profitable than honest mining. Next, we connected the attack to double-spending, as $L>k$ implies double-spending. Thus, given $\alpha$ and $\gamma$, the maximum profitable stubbornness values determine the values of $k$ where $k$-confirmation rule is at risk of double-spending which comes at no-cost to the adversary. We also provided an exact formula for the double-spending probabilities (i.e., the risk) associated with each attack cycle. We provided the minimum double-spend value needed for an attack to be profitable in the regimes where the scheme is less profitable than honest mining. In order to conceal the double-spending attack as well as to increase the associated successful double-spending probabilities, we modified the attack that does not match every honest block during the stubborn phase. We provided all the relevant analysis for this modified attack as well.
There are two possible extensions to our paper that can further increase the action space and combining the two would eventually allow all possible state-space considered in [6]. First, note that, we define the attack cycles in a way such that they are i.i.d. This is a design choice which allows us to know whether $tx$ is confirmed or not and use Doob’s optional stopping theorem in our analysis as well as the techniques provided in Appendix A and Appendix B. In this paper, each time an attack cycle starts, we have $A=0$ and $H=0$, since we essentially keep track of the length of the $A$-chain, even if the $H$-chain has some prefix consisting of adversarial blocks of the $A$-chain. On the other hand, one could consider the situations where the duration of attack cycles are defined with respect to the change in the prefix of the $H$-chain, i.e., restricting $H$-chain to consist of the honest blocks exclusively, which is essentially considered in the MDP defined in [6]. This change not only complicates the rigorous analysis and requires us to keep a parameter that tracks the number of honest blocks, but also requires us to keep track of the depth of the $tx$ and $tx^{\prime}$, hence our design choice is to avoid this. The other extension is more obvious, which is to allow trail stubbornness, i.e., to allow the adversary to continue the attack even if it falls behind. By considering an extra parameter for the upper bound on the deficit of the $A$-chain compared to the $H$-chain, our analysis can be redone, however, this would also complicate the analysis provided in Appendix A and Appendix B.
Appendix A Appendix A: Proof of Theorem 3 and Algorithm 1
Let $x_{i}$ and $y_{i}$ be the number of adversarial and honest blocks that make it into the offset chain at the end of an attack cycle $i$. Clearly, the revenue in the long-run is
$\rho=\frac{\sum_{i=1}^{T}x_{i}}{{\sum_{i=1}^{T}x_{i}+\sum_{i=1}^{T}y_{i}}},$ (43)
where $T$ is large. Moving terms, we get $\liminf_{T}\frac{1}{T}\sum_{i}\left((1-\rho)x_{i}-\rho y_{i}\right)=0$, which is the linearization model of MDP in [6]. Next, we show that it can be narrowed down to a specific case of a single cycle, where a decision about overwrite has to be made or delayed, hence we only consider $w_{\rho}(x,y)=(1-\rho)x-\rho y$ and assume $A=H+1=L$.
Let $w^{(L)}_{\rho}(x,y)$ denote the average reward associated with the current cycle of the $L$-stubborn mining strategy when $A=x$ and $H=y$ (cycle has not ended yet necessarily). For example, when $A\geq L$ and $H=A-1$, $w^{(L)}_{\rho}(A,A-1)=(1-\rho)A$ since the decision is to release the $A$-chain immediately. On the other hand, if $A+1=H\leq L$, then $w^{(L)}_{\rho}(H-1,H)=-\rho H$ since the decision is to accept the $H$-chain and abort the attack cycle.
Lemma 3
If $\rho_{L+1}<\rho_{L}$, then
$w^{(L+1)}_{\rho_{L}}(L,L-1)<w^{(L+1)}_{\rho_{L+1}}(L,L-1)<w^{(L)}_{\rho_{L}}(L,L-1)<w^{(L)}_{\rho_{L+1}}(L,L-1).$ (44)
Proof:  Let $q_{L}$ denote the probability that a cycle passes through $A=H+1=L$. Then,
00 0 $=\liminf_{T}\frac{1}{T}\sum_{i=1}^{T}w^{(L)}_{\rho_{L}}(x_{i},y_{i})$ (45)
$=q_{L}w^{(L)}_{\rho_{L}}(L,L-1)+(1-q_{L})w^{(L)}_{\rho_{L}}(\overline{L,L-1}),$ (46)
where $w^{(L)}_{\rho_{L}}(\overline{L,L-1})$ is the average reward of $L$-stubborn mining strategy associated with $(\overline{L,L-1})$, i.e., all cycles that do not pass through $A=H+1=L$ Next, for any $\rho\in[0,1]$
$w^{(L)}_{\rho}(\overline{L,L-1})=w^{(L+1)}_{\rho}(\overline{L,L-1}),$ (47)
which is due to the fact that both strategies have the same rewards for all possible arrival paths that do not pass through the event $A=H+1=L$ during a cycle. To see this, consider $4$-stubborn mining and $5$-stubborn mining, for example. All red crosses represented in Figure 1 for $4$-stubborn mining are also present in $5$-stubborn mining, i.e., the cycles have the same decisions as long as $A<4$. Similarly, all blue crosses except the blue cross at $(4,3)$ represented in Figure 1 for $4$-stubborn mining are also present in $5$-stubborn mining, i.e., the cycles have the same decisions as long as $A>4$. Thus, the only difference in rewards is, if a cycle passes through the point $(4,3)$.
Further, if $\rho_{L+1}<\rho_{L}$, as $w_{\rho}$ is decreasing in $\rho$, we have
$w^{(L)}_{\rho_{L}}(\overline{L,L-1})<w^{(L+1)}_{\rho_{L+1}}(\overline{L,L-1}).$ (48)
Thus, rewriting (46) for $(L+1)$-stubborn mining strategy
00 0 $=q_{L}w^{(L+1)}_{\rho_{L+1}}(L,L-1)+(1-q_{L})w^{(L+1)}_{\rho_{L+1}}(\overline{L,L-1})$ (49)
$>q_{L}w^{(L+1)}_{\rho_{L+1}}(L,L-1)+(1-q_{L})w^{(L)}_{\rho_{L}}(\overline{L,L-1})$ (50)
$=q_{L}\left(w^{(L+1)}_{\rho_{L+1}}(L,L-1)-w^{(L)}_{\rho_{L}}(L,L-1)\right),$ (51)
where (50) follows from (48) and (51) follows from (46). This proves the inequality in the middle of (44). The other two inequalities follow from the fact that $w_{\rho}$ is decreasing in $\rho$. $\blacksquare$
Corollary 1
If $\rho_{L+1}>\rho_{L}$, then
$w^{(L+1)}_{\rho_{L}}(L,L-1)>w^{(L+1)}_{\rho_{L+1}}(L,L-1)>w^{(L)}_{\rho_{L}}(L,L-1)>w^{(L)}_{\rho_{L+1}}(L,L-1).$ (52)
Proof:  Simply change the direction of the inequalities from (48) onwards in the proof of Lemma 3. $\blacksquare$
Lemma 4
Given $A=H+1=L$, the average result associated with $(L+1)$-stubborn mining strategy is related to $L$-stubborn mining strategy in the following way
$w^{(L+1)}_{\rho}(L,L-1)=w^{(L)}_{\rho}(L,L-1)+\beta^{2}\left(\frac{1-{\rho}}{1-2\alpha}-\frac{1}{\gamma}+\frac{(1-\gamma)^{L+1}}{\gamma}\right).$ (53)
Proof:  Clearly, if $L$-stubborn mining strategy is followed, we get $w^{(L)}_{\rho}(L,L-1)=w_{\rho}(L,0)$. On the other hand, if $(L+1)$-stubborn mining strategy is followed, the decision is delayed and $w^{(L+1)}_{\rho}(L,L-1)$ depends on the arrival of the next blocks in the following way:
  1. 1.
    The next block is adversarial w.p. $\alpha$. In this case, the adversarial advantage $A-H=2$ implies that the decision will be made when the advantage reduces to $1$, which happens after $\frac{1}{1-2\alpha}$ block arrivals in expectation and $\frac{\alpha}{1-2\alpha}$ of those blocks will be adversarial. Hence, w.p. $\alpha$ we get $w_{\rho}(L+1+\frac{\alpha}{1-2\alpha},0)$.
  2. 2.
    The next block is honest followed by an adversarial block w.p. $\beta\alpha$. As a result, we get $w_{\rho}(L+1,0)$ w.p. $\beta\alpha$.
  3. 3.
    The next two blocks are honest w.p. $\beta^{2}$. In this situation, $w_{\rho}$ depends on how many adversarial blocks are contained in the prefix of the honest block at height $L+1$:
    1. (a)
      The prefix contains $L-i>0$ adversarial blocks w.p. $\gamma(1-\gamma)^{i}$ and we have $w_{\rho}(L-i,i+1)$.
    2. (b)
      The prefix does not contain any adversarial block w.p. $(1-\gamma)^{L}$ and we have $w_{\rho}(0,L+1)$.
Combining the cases above in expectation, for $(L+1)$-stubborn mining strategy, we obtain
$w^{(L+1)}_{\rho}(L,L-1)=$ $\alpha w_{\rho}(L+1+\frac{\alpha}{1-2\alpha},0)+\beta\alpha w_{\rho}(L+1,0)+\beta^{2}(1-\gamma)^{L}w_{\rho}(0,L+1)$
$+\sum_{i=0}^{L-1}\beta^{2}\gamma(1-\gamma)^{i}w_{\rho}(L-i,i+1),$ (54)
which results in (53). $\blacksquare$
The intermediary results above are now sufficient to prove Theorem 3.
Let $u$ and $v_{a}$ be,
$u=(1-\gamma),\quad\quad v_{a}=\frac{1}{u}\left(1-\gamma\frac{1-\rho_{a}}{1-2\alpha}\right).$ (55)
If $\rho_{L}\geq\rho_{L+1}$, picking $\rho=\rho_{L}$ and $\rho=\rho_{L+1}$ in (53) implies
$v_{L}\geq v_{L+1}$ $\geq u^{L},$ (56)
due to Lemma 3.
Next, assume $\rho_{L+1}<\rho_{L+2}$ for the sake of contradiction. Rewriting (53) and (52) for the relation between $(L+1)$ and $(L+2)$-stubborn mining and picking $\rho=\rho_{L+1}$ in (53), we get
$u^{L+1}>v_{L+1},$ (57)
which creates a contradiction with (56) since $u^{L}>u^{L+1}$. The rest is just applying the recursion and the same proof applies for both claims in the theorem statement.
Proposition 1
Algorithm 1 outputs $L^{*}$.
Proof:  First, assume $L^{*}=\infty$, then clearly $L^{(1)}=\infty$ and $\rho_{L+1}>\rho_{L}$ implies
$u^{L}>v_{L},$ (58)
for all $L$. Taking limit of both sides implies $0\geq v_{L^{(1)}}$, thus, algorithm returns $\infty$.
Next, assume $L^{*}$ is finite. Then, for all $L>L^{*}$, $\rho_{L}>\rho_{L+1}$, which implies
$u^{L}<v_{L}.$ (59)
Taking limit of both sides further implies that
$0<v_{\infty}\leq\max\{v_{1},v_{2},v_{\infty}\},$ (60)
i.e., the algorithm does not return infinity. Next assume $L^{*}=1$, then clearly $L^{(1)}=1$ and the algorithm returns $1$.
If $1<L^{*}<\infty$, then clearly $L^{(1)}=2$ or $L^{(1)}=\infty$, and the algorithm enters the while loop. From Lemma 3, Lemma 4 and Corollary 1, if $\rho_{L}>\rho_{L+1}$
$\frac{\log v_{L}}{\log u}<\frac{\log v_{L+1}}{\log u}<L,$ (61)
and if $\rho_{L}>\rho_{L-1}$
$\frac{\log v_{L-1}}{\log u}>\frac{\log v_{L}}{\log u}>L-1.$ (62)
Further, note that $v_{L^{*}}\geq v_{a}$ for any $a\in\mathbb{Z}$. Assuming $L^{*}$ is finite, (62) implies, for all $i$
$L^{(i+1)}=\left\lceil{\frac{\log v_{L^{(i)}}}{\log u}}\right\rceil\geq L^{*}.$ (63)
If $L^{(i)}>L^{*}$, then, (61) implies
$L^{(i+1)}=\left\lceil{\frac{\log v_{L^{(i)}}}{\log u}}\right\rceil<L^{(i)}-1.$ (64)
and by
$L^{*}\geq\frac{\log v_{L^{*}}}{\log u}>L^{*}-1,$ (65)
the result $L^{(i)}$ converges in at most $L^{(2)}-L^{*}$ loops to $L^{*}$. It remains to show $L^{(2)}<\infty$ when $L^{*}$ is finite. By (61), we have
$L^{(2)}=\min\left\{\left\lceil\frac{\log v_{2}}{\log u}\right\rceil,\left\lceil\frac{\log v_{\infty}}{\log u}\right\rceil\right\}\leq\left\lceil\frac{\log v_{\infty}}{\log u}\right\rceil<\infty,$ (66)
which completes the proof. Note that, if there are multiple $L$ with optimal value, then the algorithm outputs the largest of those $L$. $\blacksquare$
We remark here that for the results of Section 6, we use $10^{-4}$ precision for $\alpha$ and $\gamma$ values and Algorithm 1 runs the while loop at most $4$ times.
Appendix B Appendix B: Proof of Theorem 7 and Algorithm 2
Let $w^{(S)}_{\sigma}(x,y)$ denote the average reward associated with the current cycle of the $S$-stealth mining strategy when $A=x$ and $H=y$.
Lemma 5
If $\sigma_{S+1}<\sigma_{S}$, then
$w^{(S+1)}_{\sigma_{S}}(S-1,S)+\frac{2(2S-1)}{S+1}\frac{\alpha}{1-\alpha}w^{(S+1)}_{\sigma_{S}}(S,S-1)$
$<w^{(S+1)}_{\sigma_{S+1}}(S-1,S)+\frac{2(2S-1)}{S+1}\frac{\alpha}{1-\alpha}w^{(S+1)}_{\sigma_{S+1}}(S,S-1)$
$<w^{(S)}_{\sigma_{S}}(S-1,S)+\frac{2(2S-1)}{S+1}\frac{\alpha}{1-\alpha}w^{(S)}_{\sigma_{S}}(S,S-1)$
$<w^{(S)}_{\sigma_{S+1}}(S-1,S)+\frac{2(2S-1)}{S+1}\frac{\alpha}{1-\alpha}w^{(S)}_{\sigma_{S+1}}(S,S-1).$ (67)
Similarly, if $\sigma_{S+1}>\sigma_{S}$, we can switch the direction of the inequalities above.
Proof:  As usual, $q_{S}$ denotes the probability that a cycle passes through $A=H+1=S$. Further, let $p_{S}$ denote the probability that a cycle passes through $H=A+1=S$. Then,
00 0 $=\liminf_{T}\frac{1}{T}\sum_{i=1}^{T}w^{(S)}_{\sigma_{S}}(x_{i},y_{i})$ (68)
$=q_{S}w^{(S)}_{\sigma_{S}}(S,S-1)+p_{S}w^{(S)}_{\sigma_{S}}(S-1,S)+(1-q_{S}-p_{S})w^{(S)}_{\sigma_{S}}(\overline{S}),$ (69)
where $w^{(S)}_{\sigma_{S}}(\overline{S})$ is the average reward of $S$-stealth mining strategy associated with $\overline{S}$, i.e., all cycles that do not pass through $A=H+1=S$ or $H=A+1=S$. Next, for any $\sigma\in[0,1]$
$w^{(S)}_{\sigma}(\overline{S})=w^{(S+1)}_{\sigma}(\overline{S}),$ (70)
which is due to the fact that both strategies have the same rewards for all possible arrival paths that do not pass through the event $A=H+1=S$ or $H=A+1=S$ during a cycle. Here, unlike $L$-stubborn mining, we need to exclude $H=A+1=S$ as well. To see this, notice that $(S+1)$-stealth mining does not match the honest block at height $S-1$, whereas $S$-stealth mining does, hence their rewards differ and we cannot apply the separation of dimensions argument explained in Figure 1.
Further, if $\sigma_{S+1}<\sigma_{S}$, then
$w^{(S)}_{\sigma_{S}}(\overline{S})<w^{(S+1)}_{\sigma_{S+1}}(\overline{S}).$ (71)
Thus, rewriting (69) for $(S+1)$-stealth mining strategy
00 0 $=q_{S}w^{(S+1)}_{\sigma_{S+1}}(S,S-1)+p_{S}w^{(S+1)}_{\sigma_{S+1}}(S-1,S)+(1-q_{S}-p_{S})w^{(S+1)}_{\sigma_{S+1}}(\overline{S})$ (72)
$>q_{S}w^{(S+1)}_{\sigma_{S+1}}(S,S-1)+p_{S}w^{(S+1)}_{\sigma_{S+1}}(S-1,S)+(1-q_{S}-p_{S})w^{(S)}_{\sigma_{S}}(\overline{S})$ (73)
$=q_{S}\left(w^{(S+1)}_{\sigma_{S+1}}(S,S-1)-w^{(S)}_{\sigma_{S}}(S,S-1)\right)+p_{S}\left(w^{(S+1)}_{\sigma_{S+1}}(S-1,S)-w^{(S)}_{\sigma_{S}}(S-1,S)\right).$ (74)
Notice, a cycle passes through $H=A+1=S$ w.p.
$p_{S}$ $=P[S-1,S-1]\alpha^{S-1}\beta^{S}$ (75)
$=C_{S-1}\alpha^{S-1}\beta^{S},$ (76)
and through $A=H+1=S$ w.p.
$q_{S}$ $=P[S,S-1]\alpha^{S}\beta^{S-1}$ (77)
$=C_{S}\alpha^{S}\beta^{S-1}$ (78)
$=\frac{2(2S-1)}{S+1}\cdot\frac{\alpha}{1-\alpha}\cdot p_{S},$ (79)
which concludes the proof. $\blacksquare$
Lemma 6
Given $A=H+1=S$, the average result associated with $(S+1)$-stealth mining strategy is related to $S$-stealth mining strategy in the following way,
$w^{(S+1)}_{\sigma}(S,S-1)=w^{(S)}_{\sigma}(S,S-1)-\beta^{2}\left((S+1)-\gamma S-\frac{1-\sigma}{1-2\alpha}\right).$ (80)
Proof:  Clearly, if $S$-stealth mining strategy is followed, we get $w^{(S)}_{\sigma}(S,S-1)=w_{\sigma}(S,0)$. On the other hand, if $(S+1)$-stealth mining strategy is followed, the decision is delayed and $w^{(S+1)}_{\sigma}(S,S-1)$ depends on the arrival of the next blocks in the following way:
  1. 1.
    The next block is adversarial w.p. $\alpha$. In this case, the adversarial advantage $A-H=2$ implies that the decision will be made when the advantage reduces to $1$, hence w.p. $\alpha$ we get $w_{\sigma}(S+1+\frac{\alpha}{1-2\alpha},0)$.
  2. 2.
    The next block is honest followed by an adversarial block w.p. $\beta\alpha$. As a result, we get $w_{\sigma}(S+1,0)$ w.p. $\beta\alpha$.
  3. 3.
    The next two blocks are honest w.p. $\beta^{2}$. In this situation, $w_{\sigma}$ depends on whether the last honest block switches prefix or not:
    1. (a)
      The prefix switches to adversarial blocks w.p. $\gamma$ and we have $w_{\sigma}(S,1)$.
    2. (b)
      The prefix does not contain any adversarial block w.p. $(1-\gamma)$ and we have $w_{\sigma}(0,S+1)$.
Combining the cases above in expectation, for $(S+1)$-stealth mining strategy, we obtain
$w^{(S+1)}_{\sigma}(S,S-1)=$ $\alpha w_{\sigma}(S+1+\frac{\alpha}{1-2\alpha},0)+\beta\alpha w_{\sigma}(S+1,0)+\beta^{2}\gamma w_{\sigma}(S,1)$
$+\beta^{2}(1-\gamma)w_{\sigma}(0,S+1),$ (81)
which results in (80). $\blacksquare$
Lemma 7
Given $A-1=H=S$, the average result associated with $(S+1)$-stealth mining strategy is related to $S$-stealth mining strategy in the following way,
$w^{(S+1)}_{\sigma}(S-1,S)=w^{(S)}_{\sigma}(S-1,S)-\gamma(S-1).$ (82)
Proof:  Clearly, if $(S+1)$-stealth mining strategy is followed, we get $w^{(S+1)}_{\sigma}(S-1,S)=w_{\sigma}(0,S)=-\sigma S$. On the other hand, if $S$-stealth mining strategy is followed, the reward depends on the prefix of the last honest block:
  1. 1.
    The prefix contains no adversarial blocks w.p. $1-\gamma$. In this case, we get $w_{\sigma}(0,S)$.
  2. 2.
    The prefix contains $S-1$ adversarial blocks w.p. $\gamma$. In this case, we get $w_{\sigma}(S-1,1)$.
Combining the cases above results in (82). $\blacksquare$
The intermediary results above are now sufficient to prove Theorem 7.
If $\sigma_{S}\geq\sigma_{S+1}$, picking $\sigma=\sigma_{S+1}$ in (80) and (82) implies
$\gamma\frac{S^{2}-1}{2(2S-1)}+\alpha\beta(S+1-\gamma S)\geq\alpha\beta\frac{1-\sigma_{S+1}}{1-2\alpha},$ (83)
due to Lemma 5.
If $\sigma_{S+1}<\sigma_{S+2}$, rewriting Lemma 5 for the relation between $(S+1)$ and $(S+2)$-stealth mining and picking $\sigma=\sigma_{S+1}$ in (80) and (82), we get
$\alpha\beta\frac{1-\sigma_{S+1}}{1-2\alpha}>\gamma\frac{S(S+2)}{2(2S+1)}+\alpha\beta(S+2-\gamma(S+1)).$ (84)
Combining the two implies,
$0>\gamma\frac{2S^{2}+1}{4S^{2}-1}+2\alpha\beta(1-\gamma),$ (85)
which creates a contradiction since the right hand side is nonnegative for $S\geq 1$, $0.5\geq\alpha\geq 0$ and $1\geq\gamma\geq 0$. This concludes the proof of Theorem 7.
It is clear that $f(\sigma)<\infty$ for $\sigma\in[0,1]$. Further, one can also show that
$S^{*}\geq f(\sigma_{S^{*}})>S^{*}-1,$ (86)
by Theorem 7. This in turn, by modifying the arguments of Proposition 1 with the results proven in this section, implies that Algorithm 2 converges to $S^{*}$.